population with a mean of 110 and a standard deviation of 6, what is the likelihood of a score of 100 or lower?
Answer by stanbon(75887)

population with a mean of 110 and a standard deviation of 6, what is the likelihood of a score of 100 or lower?
------
z(100) = (100-110)/6 = -10/6 = -5/3
--------
P(x <=100) = P(z <= -5/3) = normalcdf(-100,-5/3) = 0.4779
--------------
Cheers,
Stan H.
==============